OpenStudy (ggalday000):

Given f(x) = 3x – 5, what is f(–4)? A. B. C. D.

OpenStudy (ggalday000):

A is 1/3 B is -6 C is -17 D is 7

Mehek (mehek14):

Plug -4 instead of x f(-4) = 3*-4 - 5

rebeccaxhawaii (rebeccaxhawaii):

^^

Mehek (mehek14):

Start with multiplication 3 * (-4) = ?

geerky42 (geerky42):

\[f(-4)=3(-4)-5\\\phantom{f(-4)}=-12-5\\\phantom{f(-4)}=\boxed{-17}\]

OpenStudy (ggalday000):

ohhh ok thank you

Mehek (mehek14):

Yw :)